# From Pixels to Prose: Advancing Multi-Modal Language Models for Remote Sensing

### v1

@@ Abstract
Index Terms:
Multi-Modal Language Models, Remote Sensing, Self-Supervised Learning, Cross-Modal Fusion, Transformer Architectures
@@ I Introduction
Remote sensing has evolved from basic image capture to advanced systems that integrate visual and textual data processing. This progression has led to the development of multi-modal language models (MLLMs) capable of interpreting and describing satellite imagery using natural language [1, 2]. Such integration enhances the analysis of Earth observation data, making it more accessible and informative.
Applying MLLMs to remote sensing data has significantly improved automated Earth observation analysis. These models support vital applications in environmental monitoring, urban planning, and disaster response by enabling more efficient and accurate extraction of information from satellite imagery [3, 4]. This advancement enhances decision-making processes across various sectors.
@@
Remote Sensing, root, [Future Directions
and Conclusions, level1, [Potential Applications
[3, 5, 1, 6, 7, 8], leaf] [Research Opportunities
[3, 1], leaf] ] [Challenges and
Limitations, level1, [Application-Specific
Challenges, level2, [Domain Adaptation [9, 4, 10, 11], leaf] ] [Technical Challenges, level2, [Computational Requirements [12, 13, 14], leaf] [Model Scalability [15, 16, 17], leaf] [Data Quality Issues [18, 19, 20, 11], leaf] ] ] [Datasets and
Resources, level1, [Evaluation Metrics, level2, [CIDEr [21], leaf] [METEOR [22], leaf] [BLEU [23], leaf] ] [Training Resources, level2, [Training Frameworks [24, 25], leaf] [Pre-trained Models [3], leaf] ] [Benchmark Datasets, level2, [ChatEarthNet [26], leaf] [RS5M [9], leaf] [RSICap [27], leaf] ] ] [Current Applications
and
Implementations, level1, [Cross-Modal
Applications, level2, [Visual Question Answering [28, 29, 30, 31], leaf] [Image-to-Text Generation [32, 33, 34, 35], leaf] [Text-to-Image Retrieval [36, 4, 10, 11], leaf] ] [Image Understanding, level2, [Change Detection [37, 38, 39], leaf] [Object Detection [40, 41, 42, 43, 44, 45, 46], leaf] [Scene Description [47, 34, 48, 49], leaf] ] ] [Technical
Foundations, level1, [Remote Sensing Data
Characteristics, level2, [Temporal Aspects [37, 50, 51, 52], leaf] [Spectral Information [53, 54, 55, 56], leaf] [Spatial Resolution [3, 57, 58, 59], leaf] ] [Multi-Modal
Language Models, level2, [Integration Approaches [60, 61, 62, 63], leaf] [Learning Mechanisms [3, 57, 64, 65], leaf] [Architecture Components [66, 67, 68], leaf] ] ] ]
@@ II-A1 Architecture Components
MLLMs in remote sensing typically employ a dual-encoder architecture, integrating specialized components for processing both visual and textual information. These models often utilize a Transformer-based structure, which has proven effective in handling the complexities of remote sensing data [66].
The visual component of these architectures commonly incorporates a Vision Transformer (ViT) or a convolutional neural network as the image encoder. This encoder is responsible for extracting relevant features from satellite imagery and other remote sensing data sources. The ViT architecture, in particular, has shown promise in adapting its attention mechanism and Transformer Encoder for remote sensing applications, allowing for efficient processing of spatial information [67].
On the language side, these models frequently employ BERT or similar Transformer-based language models as the text encoder. This component is crucial for understanding and processing textual queries, captions, or metadata associated with remote sensing imagery. The language encoder’s self-attention mechanisms enable it to capture complex linguistic relationships and context [68].
A key aspect of MLLMs in remote sensing is the integration of these visual and textual components. This integration is often achieved through cross-modal fusion techniques and sophisticated attention mechanisms. For instance, some architectures use self-attention multi-modal encoders that combine convolutional and recurrent elements with cross-modal fusion components [66]. Others employ unified Transformer models with cross-modal mixture experts, which allow for more nuanced interactions between visual and textual features [69].
The attention mechanism plays a crucial role in these architectures, enabling the model to focus on relevant parts of both the image and text inputs. This is particularly important in remote sensing applications, where the ability to identify and focus on specific geographical features or phenomena is essential. The self-attention layers in these models allow for dynamic weighting of different input elements, enhancing the model’s capacity to understand complex spatial and semantic relationships [68].
@@ II-A2 Learning Mechanisms
Self-supervised learning techniques form the foundation of modern multi-modal models in remote sensing, enabling them to leverage vast amounts of unlabeled satellite imagery and associated text descriptions. These techniques allow models to learn meaningful representations without relying on expensive and time-consuming manual annotations. For instance, the SkySense model employs a Multi-Granularity Contrastive Learning approach for self-supervised pre-training on a large-scale multi-modal remote sensing imagery dataset, demonstrating the effectiveness of this learning mechanism in capturing complex spatiotemporal relationships [3].
Contrastive learning has emerged as a powerful paradigm in developing robust multi-modal representations for remote sensing applications. This approach helps establish strong relationships between visual and textual representations, enhancing the model’s ability to align information across modalities. For example, Zhu et al. propose a cross-modal contrastive learning method that incorporates spatio-temporal context for multi-scale remote sensing image retrieval. Their approach uses a self-supervised contrastive loss to improve the model’s capacity to capture correlations between different modalities and scales [57].
Cross-modal training is crucial in bridging the gap between different data modalities in remote sensing, such as optical imagery, synthetic aperture radar (SAR) data, and textual information. Zavras et al. address this challenge by developing a methodology for aligning distinct remote sensing imagery modalities with visual and textual modalities of CLIP (Contrastive Language-Image Pre-training). Their two-stage procedure involves robust fine-tuning of CLIP to handle distribution shifts, followed by cross-modal alignment of a remote sensing modality encoder [64].
The integration of intra-modal and cross-modal learning strategies has shown promise in improving the overall performance of multi-modal models. Wu et al. demonstrate this by combining self-supervised intra-modal and cross-modal contrastive learning for point cloud understanding, which is applicable to remote sensing data. Their approach leverages the complementary nature of different modalities to enhance feature representations and improve downstream task performance [65].
@@ II-A3 Integration Approaches
Cross-modal feature fusion techniques play a crucial role in combining visual and textual representations in MLLMs for remote sensing applications. These techniques often employ carefully designed attention mechanisms and alignment strategies to ensure effective information exchange between modalities while preserving the unique characteristics of each data type. For instance, Liu and Wang propose a bidirectional feature fusion approach that incorporates enhanced alignment for multimodal semantic segmentation of remote sensing images. Their method uses cross-attention mechanisms to maintain the integrity of single-modal information while achieving effective feature alignment across modalities [60].
Modal alignment is a critical aspect of integration approaches, as it ensures that features from different modalities are properly synchronized and correlated. Sun et al. introduce a cross-modal pre-aligned method that incorporates both global and local information for remote sensing image and text retrieval. Their approach employs a multi-modal encoder for cross-modal feature fusion and utilizes cross-attention mechanisms to fully exploit the relationships between different modalities. This pre-alignment strategy helps to establish a strong foundation for subsequent feature integration and retrieval tasks [61].
Cross-attention mechanisms have emerged as a powerful tool for facilitating information exchange and alignment between different modalities in remote sensing applications. Ma et al. present a crossmodal multiscale fusion network for semantic segmentation of remote sensing data, which leverages cross-attention and cross-transformer techniques for multimodal data fusion. Their approach addresses challenges in feature extraction, alignment, and fusion across different modalities, demonstrating the effectiveness of cross-attention in capturing complex inter-modal relationships [62].
To further enhance the integration of multi-modal remote sensing data, researchers have explored self-supervised learning techniques combined with asymmetric attention fusion. Xu et al. propose a method that utilizes cross-attention mechanisms to transfer features between modalities during pre-training. This approach addresses the challenge of integrating multi-modal data with strong inter-modal complementarity, allowing the model to learn more robust and generalizable representations [63].
@@ II-B1 Spatial Resolution Considerations
Spatial resolution in remote sensing imagery varies significantly, ranging from sub-meter to kilometer-scale ground sampling distances. This variability presents unique challenges for MLLMs, necessitating adaptive processing approaches to maintain model performance across different scales. The SkySense model, for instance, addresses this issue by employing a factorized multi-modal spatiotemporal encoder that learns representations across different modal and spatial granularities. This approach enables the model to effectively process and interpret remote sensing data at various spatial resolutions, demonstrating the importance of multi-scale analysis in handling diverse Earth observation imagery [3].
To tackle those challenges, researchers have developed multi-scale analysis techniques that allow models to adapt to different levels of detail in remote sensing imagery. Zhu et al. propose a cross-modal contrastive learning approach that incorporates spatio-temporal context for correlation-aware multi-scale remote sensing image retrieval. Their method addresses the spatial resolution variability by considering multiple scales during the learning process, enabling more robust and adaptive representations of remote sensing data [57].
The integration of multi-scale alignment methods has shown promise in improving the performance of MLLMs across different spatial resolutions. Yang et al. introduce a Multi-Scale Alignment (MSA) method for remote sensing image-text retrieval, which includes a Multi-scale Cross-Modal Alignment Transformer (MSCMAT). This approach computes cross-attention between single-scale image features and localized text features while integrating global textual context. By considering multiple scales in the alignment process, the model can better handle the diverse spatial resolutions encountered in remote sensing applications [58].
For high-resolution remote sensing imagery, adaptive processing techniques are particularly crucial. Zhang et al. present SegCLIP, a multimodal approach for high-resolution remote sensing semantic segmentation that incorporates a multi-scale prototype transformer decoder. This method specifically addresses the challenges associated with processing high-resolution remote sensing imagery by employing multi-scale analysis techniques. By adapting to different levels of detail within the high-resolution data, SegCLIP demonstrates the effectiveness of multi-scale approaches in maintaining model performance across varying spatial resolutions [59].
@@ II-B2 Spectral Information
Remote sensing systems capture data across multiple spectral bands, ranging from visible light to infrared wavelengths, providing rich information about Earth’s surface properties. This spectral diversity is a key component in developing advanced MLLMs for remote sensing applications. Multispectral (MS) sensors measure reflected light within specific wavelength ranges, offering a comprehensive view of the Earth’s surface that goes beyond what is visible to the human eye [53]. These sensors typically capture data in several discrete bands, each sensitive to a particular portion of the electromagnetic spectrum.
Hyperspectral sensors, on the other hand, collect data across a much larger number of narrower spectral bands, often numbering in the hundreds. This high spectral resolution allows for more detailed analysis of surface properties and materials. The integration of hyperspectral data into MLLMs presents both opportunities and challenges, as it provides a wealth of information but also increases computational complexity [54].
The selection and processing of spectral bands play a crucial role in extracting meaningful features for multi-modal analysis. Band selection techniques are employed to identify the most informative spectral bands for specific applications, reducing data dimensionality while preserving essential information. This process is particularly important when dealing with hyperspectral data, where the high number of bands can lead to the ”curse of dimensionality” and increased processing time [55].
Multi-modal fusion approaches, such as those employing transformer architectures, have shown promise in effectively combining spectral information with other data modalities like synthetic aperture radar (SAR) or textual data. These fusion techniques allow models to leverage the complementary nature of different spectral ranges and data types, leading to improved performance in tasks such as image classification, object detection, and semantic segmentation [56].
The development of foundation models for remote sensing has further advanced the utilization of spectral information. These models, inspired by large language models, aim to learn generalizable representations from diverse spectral data sources. By incorporating varying spatial and spectral resolutions, foundation models can potentially address the challenges posed by the complexity of Earth’s environments and the distinctive feature patterns found in remote sensing data [56].
@@ II-B3 Temporal Aspects
One significant application of temporal aspects in multi-modal models is change detection. ChangeCLIP, for instance, utilizes multimodal vision-language representation learning to identify changes in satellite imagery over time [37]. This approach demonstrates the potential of integrating temporal information with both visual and textual data to improve the accuracy and interpretability of change detection tasks in remote sensing.
The analysis of satellite image time series presents unique challenges due to the irregular nature of data acquisition. To address this, researchers have developed self-supervised learning techniques that can handle the complexities of temporal data in remote sensing [50]. These methods aim to extract meaningful temporal patterns and changes from satellite imagery, even when dealing with non-uniform sampling intervals or missing data points.
Advancements in generative models have also contributed to the field of temporal remote sensing analysis. Changen2, a multi-temporal remote sensing generative change foundation model, works with time series of remote sensing images to generate globally distributed change detection datasets [51]. This approach not only aids in change detection but also has the potential to augment existing datasets, addressing the scarcity of labeled temporal data in remote sensing.
Integrating knowledge-guided approaches with multimodal foundation models pushes the boundaries of spatio-temporal remote sensing. These models can handle a variety of tasks that require temporal understanding, such as land-use land cover change detection and crop yield prediction [52]. By incorporating domain knowledge and temporal information, these models can provide more accurate and contextually relevant interpretations of remote sensing data over time.

### v2

@@ I Introduction
Remote sensing has evolved from basic image capture to advanced systems that integrate visual and textual data processing. This progression has led to the development of multi-modal language models (MLLMs) capable of interpreting and describing satellite imagery using natural language [zhang2024earthgpt, zhan2024skyeyegpt]. Such integration enhances the analysis of Earth observation data, making it more accessible and informative.
Applying MLLMs to remote sensing data has significantly improved automated Earth observation analysis. These models support vital applications in environmental monitoring, urban planning, and disaster response by enabling more efficient and accurate extraction of information from satellite imagery [guo2024skysense, mi2024knowledge]. This advancement enhances decision-making processes across various sectors.
@@
[…]
MLLM and Remote Sensing Future Directions and Conclusions Potential Applications [guo2024skysense, yuan2024surveillance, zhang2024earthgpt, zheng2023marinegpt, peng2024securing, peng2024jailbreaking] Research Opportunities [guo2024skysense, zhang2024earthgpt] Challenges and Limitations Application-Specific Challenges Domain Adaptation [zhang2023rs5m, mi2024knowledge, pan2024pir, mikriukov2022deep] Technical Challenges Computational Requirements [basu2024understanding, hadji2023context, gopalkrishnan2024multi] Model Scalability [aghajanyan2023scaling, tian2022resformer, liu2024sphinx] Data Quality Issues [pasquarella2023comprehensive, thai2022riesz, song2020unsupervised, mikriukov2022deep] Datasets and Resources Evaluation Metrics CIDEr [vedantam2015cider] METEOR [banerjee2005meteor] BLEU [papineni2002bleu] Training Resources Training Frameworks [huang2024language, berg2023joint] Pre-trained Models [guo2024skysense] Benchmark Datasets ChatEarthNet [yuan2024chatearthnet] RS5M [zhang2023rs5m] RSICap [hu2023rsgpt] Current Applications and Implementations Cross-Modal Applications Visual Question Answering [lobry2020rsvqa, zheng2021mutual, siebert2022multi, hackel2023lit] Image-to-Text Generation [shen2020remote, li2020multi, silva2024large, zhan2023rsvg] Text-to-Image Retrieval [yuan2022remote, mi2024knowledge, pan2024pir, mikriukov2022deep] Image Understanding Change Detection [dong2024changeclip, muhtar2024lhrsbot, ferrod2024towards] Object Detection [wagner2020u, zhang2017separate, kampffmeyer2016semantic, garg2021isdnet, zang2024contextual, wang2024visionllm, zhao2024llm] Scene Description [li2024cross, silva2024large, li2024learning, zhao2021high] Technical Foundations Remote Sensing Data Characteristics Temporal Aspects [dong2024changeclip, chen2023self, zheng2024changen2, ravirathinam2024towards] Spectral Information [roy2023multimodal, kieu2024multimodal, he2023foundation, xiao2024foundation] Spatial Resolution [guo2024skysense, zhu2024cross, yang2024transcending, zhang2024segclip] Multi-Modal Language Models Integration Approaches [liu2024bidirectional, sun2024cross, ma2022crossmodal, xu2023exploring] Learning Mechanisms [guo2024skysense, zhu2024cross, zavras2024mind, wu2023self] Architecture Components [silva2022remote, hoffmann2022multi, zhang2023fusion]
@@ II-A1 Architecture Components
MLLMs in remote sensing typically employ a dual-encoder architecture, integrating specialized components for processing both visual and textual information. These models often utilize a Transformer-based structure, which has proven effective in handling the complexities of remote sensing data [silva2022remote].
The visual component of these architectures commonly incorporates a Vision Transformer (ViT) or a convolutional neural network as the image encoder. This encoder is responsible for extracting relevant features from satellite imagery and other remote sensing data sources. The ViT architecture, in particular, has shown promise in adapting its attention mechanism and Transformer Encoder for remote sensing applications, allowing for efficient processing of spatial information [hoffmann2022multi].
On the language side, these models frequently employ BERT or similar Transformer-based language models as the text encoder. This component is crucial for understanding and processing textual queries, captions, or metadata associated with remote sensing imagery. The language encoder’s self-attention mechanisms enable it to capture complex linguistic relationships and context [zhang2023fusion].
A key aspect of MLLMs in remote sensing is the integration of these visual and textual components. This integration is often achieved through cross-modal fusion techniques and sophisticated attention mechanisms. For instance, some architectures use self-attention multi-modal encoders that combine convolutional and recurrent elements with cross-modal fusion components [silva2022remote]. Others employ unified Transformer models with cross-modal mixture experts, which allow for more nuanced interactions between visual and textual features [liu2023unified].
The attention mechanism plays a crucial role in these architectures, enabling the model to focus on relevant parts of both the image and text inputs. This is particularly important in remote sensing applications, where the ability to identify and focus on specific geographical features or phenomena is essential. The self-attention layers in these models allow for dynamic weighting of different input elements, enhancing the model’s capacity to understand complex spatial and semantic relationships [zhang2023fusion].
@@ II-A2 Learning Mechanisms
Self-supervised learning techniques form the foundation of modern multi-modal models in remote sensing, enabling them to leverage vast amounts of unlabeled satellite imagery and associated text descriptions. These techniques allow models to learn meaningful representations without relying on expensive and time-consuming manual annotations. For instance, the SkySense model employs a Multi-Granularity Contrastive Learning approach for self-supervised pre-training on a large-scale multi-modal remote sensing imagery dataset, demonstrating the effectiveness of this learning mechanism in capturing complex spatiotemporal relationships [guo2024skysense].
Contrastive learning has emerged as a powerful paradigm in developing robust multi-modal representations for remote sensing applications. This approach helps establish strong relationships between visual and textual representations, enhancing the model’s ability to align information across modalities. For example, Zhu et al. propose a cross-modal contrastive learning method that incorporates spatio-temporal context for multi-scale remote sensing image retrieval. Their approach uses a self-supervised contrastive loss to improve the model’s capacity to capture correlations between different modalities and scales [zhu2024cross].
Cross-modal training is crucial in bridging the gap between different data modalities in remote sensing, such as optical imagery, synthetic aperture radar (SAR) data, and textual information. Zavras et al. address this challenge by developing a methodology for aligning distinct remote sensing imagery modalities with visual and textual modalities of CLIP (Contrastive Language-Image Pre-training). Their two-stage procedure involves robust fine-tuning of CLIP to handle distribution shifts, followed by cross-modal alignment of a remote sensing modality encoder [zavras2024mind].
The integration of intra-modal and cross-modal learning strategies has shown promise in improving the overall performance of multi-modal models. Wu et al. demonstrate this by combining self-supervised intra-modal and cross-modal contrastive learning for point cloud understanding, which is applicable to remote sensing data. Their approach leverages the complementary nature of different modalities to enhance feature representations and improve downstream task performance [wu2023self].
@@ II-A3 Integration Approaches
Cross-modal feature fusion techniques play a crucial role in combining visual and textual representations in MLLMs for remote sensing applications. These techniques often employ carefully designed attention mechanisms and alignment strategies to ensure effective information exchange between modalities while preserving the unique characteristics of each data type. For instance, Liu and Wang propose a bidirectional feature fusion approach that incorporates enhanced alignment for multimodal semantic segmentation of remote sensing images. Their method uses cross-attention mechanisms to maintain the integrity of single-modal information while achieving effective feature alignment across modalities [liu2024bidirectional].
Modal alignment is a critical aspect of integration approaches, as it ensures that features from different modalities are properly synchronized and correlated. Sun et al. introduce a cross-modal pre-aligned method that incorporates both global and local information for remote sensing image and text retrieval. Their approach employs a multi-modal encoder for cross-modal feature fusion and utilizes cross-attention mechanisms to fully exploit the relationships between different modalities. This pre-alignment strategy helps to establish a strong foundation for subsequent feature integration and retrieval tasks [sun2024cross].
Cross-attention mechanisms have emerged as a powerful tool for facilitating information exchange and alignment between different modalities in remote sensing applications. Ma et al. present a crossmodal multiscale fusion network for semantic segmentation of remote sensing data, which leverages cross-attention and cross-transformer techniques for multimodal data fusion. Their approach addresses challenges in feature extraction, alignment, and fusion across different modalities, demonstrating the effectiveness of cross-attention in capturing complex inter-modal relationships [ma2022crossmodal].
To further enhance the integration of multi-modal remote sensing data, researchers have explored self-supervised learning techniques combined with asymmetric attention fusion. Xu et al. propose a method that utilizes cross-attention mechanisms to transfer features between modalities during pre-training. This approach addresses the challenge of integrating multi-modal data with strong inter-modal complementarity, allowing the model to learn more robust and generalizable representations [xu2023exploring].
@@ II-B1 Spatial Resolution Considerations
Spatial resolution in remote sensing imagery varies significantly, ranging from sub-meter to kilometer-scale ground sampling distances. This variability presents unique challenges for MLLMs, necessitating adaptive processing approaches to maintain model performance across different scales. The SkySense model, for instance, addresses this issue by employing a factorized multi-modal spatiotemporal encoder that learns representations across different modal and spatial granularities. This approach enables the model to effectively process and interpret remote sensing data at various spatial resolutions, demonstrating the importance of multi-scale analysis in handling diverse Earth observation imagery [guo2024skysense].
To tackle those challenges, researchers have developed multi-scale analysis techniques that allow models to adapt to different levels of detail in remote sensing imagery. Zhu et al. propose a cross-modal contrastive learning approach that incorporates spatio-temporal context for correlation-aware multi-scale remote sensing image retrieval. Their method addresses the spatial resolution variability by considering multiple scales during the learning process, enabling more robust and adaptive representations of remote sensing data [zhu2024cross].
The integration of multi-scale alignment methods has shown promise in improving the performance of MLLMs across different spatial resolutions. Yang et al. introduce a Multi-Scale Alignment (MSA) method for remote sensing image-text retrieval, which includes a Multi-scale Cross-Modal Alignment Transformer (MSCMAT). This approach computes cross-attention between single-scale image features and localized text features while integrating global textual context. By considering multiple scales in the alignment process, the model can better handle the diverse spatial resolutions encountered in remote sensing applications [yang2024transcending].
For high-resolution remote sensing imagery, adaptive processing techniques are particularly crucial. Zhang et al. present SegCLIP, a multimodal approach for high-resolution remote sensing semantic segmentation that incorporates a multi-scale prototype transformer decoder. This method specifically addresses the challenges associated with processing high-resolution remote sensing imagery by employing multi-scale analysis techniques. By adapting to different levels of detail within the high-resolution data, SegCLIP demonstrates the effectiveness of multi-scale approaches in maintaining model performance across varying spatial resolutions [zhang2024segclip].
@@ II-B2 Spectral Information
Remote sensing systems capture data across multiple spectral bands, ranging from visible light to infrared wavelengths, providing rich information about Earth’s surface properties. This spectral diversity is a key component in developing advanced MLLMs for remote sensing applications. Multispectral (MS) sensors measure reflected light within specific wavelength ranges, offering a comprehensive view of the Earth’s surface that goes beyond what is visible to the human eye [roy2023multimodal]. These sensors typically capture data in several discrete bands, each sensitive to a particular portion of the electromagnetic spectrum.
Hyperspectral sensors, on the other hand, collect data across a much larger number of narrower spectral bands, often numbering in the hundreds. This high spectral resolution allows for more detailed analysis of surface properties and materials. The integration of hyperspectral data into MLLMs presents both opportunities and challenges, as it provides a wealth of information but also increases computational complexity [kieu2024multimodal].
The selection and processing of spectral bands play a crucial role in extracting meaningful features for multi-modal analysis. Band selection techniques are employed to identify the most informative spectral bands for specific applications, reducing data dimensionality while preserving essential information. This process is particularly important when dealing with hyperspectral data, where the high number of bands can lead to the ”curse of dimensionality” and increased processing time [he2023foundation].
Multi-modal fusion approaches, such as those employing transformer architectures, have shown promise in effectively combining spectral information with other data modalities like synthetic aperture radar (SAR) or textual data. These fusion techniques allow models to leverage the complementary nature of different spectral ranges and data types, leading to improved performance in tasks such as image classification, object detection, and semantic segmentation [xiao2024foundation].
The development of foundation models for remote sensing has further advanced the utilization of spectral information. These models, inspired by large language models, aim to learn generalizable representations from diverse spectral data sources. By incorporating varying spatial and spectral resolutions, foundation models can potentially address the challenges posed by the complexity of Earth’s environments and the distinctive feature patterns found in remote sensing data [xiao2024foundation].
@@ II-B3 Temporal Aspects
One significant application of temporal aspects in multi-modal models is change detection. ChangeCLIP, for instance, utilizes multimodal vision-language representation learning to identify changes in satellite imagery over time [dong2024changeclip]. This approach demonstrates the potential of integrating temporal information with both visual and textual data to improve the accuracy and interpretability of change detection tasks in remote sensing.
The analysis of satellite image time series presents unique challenges due to the irregular nature of data acquisition. To address this, researchers have developed self-supervised learning techniques that can handle the complexities of temporal data in remote sensing [chen2023self]. These methods aim to extract meaningful temporal patterns and changes from satellite imagery, even when dealing with non-uniform sampling intervals or missing data points.
Advancements in generative models have also contributed to the field of temporal remote sensing analysis. Changen2, a multi-temporal remote sensing generative change foundation model, works with time series of remote sensing images to generate globally distributed change detection datasets [zheng2024changen2]. This approach not only aids in change detection but also has the potential to augment existing datasets, addressing the scarcity of labeled temporal data in remote sensing.
Integrating knowledge-guided approaches with multimodal foundation models pushes the boundaries of spatio-temporal remote sensing. These models can handle a variety of tasks that require temporal understanding, such as land-use land cover change detection and crop yield prediction [ravirathinam2024towards]. By incorporating domain knowledge and temporal information, these models can provide more accurate and contextually relevant interpretations of remote sensing data over time.